OpenStudy (anonymous):

Suppose y varies directly with x, and y = 8 when x = –6. What direct variation equation relates x and y? What is the value of y when x = –2?

OpenStudy (anonymous):

y=mx 8=-6m m=-4/3 y=(-4/3)*(-2)

OpenStudy (anonymous):

I am confused is that your answer

OpenStudy (anonymous):

A. y = –0.75x; 1.50 B. y = –1.33x; 2.67 C. y = 1.33x; –2.67 D. y = 0.13x; –0.25 these are the answers that are given to choose from

OpenStudy (anonymous):

m is the slope. y=m*x

OpenStudy (anonymous):

-4/3=-1.33 B